OpenStudy (anonymous):

3. 27a*0=0 (1 point) A. Associative property of multiplication B. commutative property of multiplication C. zero property of multiplication D. identity property of multiplication

OpenStudy (anonymous):

what do you think the answer is?

OpenStudy (anonymous):

A?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

an example of Associative property of multiplication is (2 * 3) * 4 = 2 * (3 * 4)

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

is it C?

OpenStudy (anonymous):

FRICK.

OpenStudy (anonymous):

B

OpenStudy (anonymous):

NO wait the answer is C. Sorry.

OpenStudy (anonymous):

An example of zero property of multiplication is 6 * 0=0, 0=0 * 6

OpenStudy (anonymous):

27a*0= 0 and 0*27a= 0

OpenStudy (anonymous):

oh ok i thought so lol